OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (5,1) and (9,-9)

OpenStudy (anonymous):

Help Please?

OpenStudy (anonymous):

slope = (-9-1) / (9-5) = -10 / 4 = -5/2

OpenStudy (anonymous):

Thanks !